# Bounding the SNPR distance between two tree-child networks using generalised agreement forests

### v1

@@ 6. Bounding the tree-child SNPR distance
Figure 4. The two cases in the construction of $R_{c_{u^{\prime}}}$ from $R^{\prime}_{c_{u^{\prime}}}$. Red indicates vertices and edges of $R_{c_{u^{\prime}}}$ and $R^{\prime}_{c_{u^{\prime}}}$.
Hence, we may assume that $e\in{\mathcal{R}}^{\prime}_{t-1}$. Let $R^{\prime}_{u}$ be the element in ${\mathcal{R}}^{\prime}_{t-1}$ that contains $e$, let $R^{\prime}_{c_{u}}$ be the element in ${\mathcal{R}}^{\prime}_{t-1}$ that contains $c_{u}$, and let $R^{\prime}_{c_{u^{\prime}}}$ be the element in ${\mathcal{R}}^{\prime}_{t-1}$ that contains $c_{u^{\prime}}$. Recall that $R^{\prime}_{u}$, $R^{\prime}_{c_{u}}$, and $R^{\prime}_{c_{u^{\prime}}}$ are not necessarily pairwise distinct. If $(p_{u^{\prime}},c_{u^{\prime}})\in R^{\prime}_{c_{u^{\prime}}}$, then set $R_{c_{u^{\prime}}}$ to be the directed graph obtained from $R^{\prime}_{c_{u^{\prime}}}$ by subdividing $(p_{u^{\prime}},c_{u^{\prime}})$ with a new vertex $u^{\prime}$. Otherwise, if $(p_{u^{\prime}},c_{u^{\prime}})\notin R^{\prime}_{c_{u^{\prime}}}$, then set $R_{c_{u^{\prime}}}$ to be the directed graph obtained from $R^{\prime}_{c_{u^{\prime}}}$ by adding $(u^{\prime},c_{u^{\prime}})$. The construction is shown in Figure 4. Intuitively, $R_{c_{u^{\prime}}}$ is an extension of a component of a phylogenetic digraph in ${\mathcal{N}}_{t}$. Lastly, if $R_{u}^{\prime}=R^{\prime}_{c_{u^{\prime}}}$, then reset $R^{\prime}_{u}=R_{c_{u^{\prime}}}$ and, if $R_{c_{u}}^{\prime}=R^{\prime}_{c_{u^{\prime}}}$, then reset $R^{\prime}_{c_{u}}=R_{c_{u^{\prime}}}$.
@@
    If $W$ contains $u$ and $|W|\geq 2$, and $R^{\prime}_{u}\neq R^{\prime}_{c_{u}}$, then obtain $R_{u}$ from $R^{\prime}_{u}$ by deleting each vertex in $W$, and obtain $R_{c_{u}}$ from $R^{\prime}_{c_{u}}$ by adding $(p_{u},c_{u})$ and each edge of $R^{\prime}_{u}$ that joins two vertices in $W-\{u\}$.
@@
    If $W$ contains $u$ and $|W|\geq 2$, and $R^{\prime}_{u}=R^{\prime}_{c_{u}}$, then obtain $R_{u}$ from $R^{\prime}_{u}$ by deleting $u$ and adding $(p_{u},c_{u})$.
As an aside, recall that we are dealing with extensions (and not with the more restricted root extensions). Indeed, if $c_{u}$ is a reticulation in ${\mathcal{N}}_{t-1}$ and a vertex with in-degree one and out-degree one in $R_{c_{u}}^{\prime}$, then $R_{c_{u}}$ is an extension and not a root extension. Now, regardless of which of (R1)–(R4) applies, let ${\mathcal{R}}^{\prime}_{t}=({\mathcal{R}}^{\prime}_{t-1}-\{R^{\prime}_{u},R^{\prime}_{c_{u}},R^{\prime}_{c_{u^{\prime}}}\})\cup\{R_{u},R_{c_{u}},R_{c_{u^{\prime}}}\}$. It is easily checked that ${\mathcal{R}}^{\prime}_{t}$ is an extension of ${\mathcal{D}}^{\prime}$ in ${\mathcal{N}}_{t}$ with
@@
$\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$$\leq$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}^{\prime}_{t}}|)$
@@
We complete the proof of the lemma by assuming that $e\in{\mathcal{M}}^{\prime}_{t-1}$. Let $e_{{\mathcal{D}}^{\prime}}=(u_{{\mathcal{D}}^{\prime}},v_{{\mathcal{D}}^{\prime}})$ be the unique edge in ${\mathcal{D}}^{\prime}$ that $e$ corresponds to. If $u_{{\mathcal{D}}^{\prime}}$ has in-degree two in ${\mathcal{D}}^{\prime}$, let $p_{{\mathcal{D}}^{\prime}}$ and $p^{\prime}_{{\mathcal{D}}^{\prime}}$ be the two parents of $u_{{\mathcal{D}}^{\prime}}$. Furthermore, if $u_{{\mathcal{D}}^{\prime}}$ has out-degree two, let $v^{\prime}_{{\mathcal{D}}^{\prime}}$ be the child of $u_{{\mathcal{D}}^{\prime}}$ that is not $v_{{\mathcal{D}}^{\prime}}$ and, if $v^{\prime}_{{\mathcal{D}}^{\prime}}$ is a reticulation, let $p^{\prime\prime}_{{\mathcal{D}}^{\prime}}$ be the parent of $v^{\prime}_{{\mathcal{D}}^{\prime}}$ that is not $u_{{\mathcal{D}}^{\prime}}$. Since ${\mathcal{D}}^{\prime}$ is tree-child, observe that each of $p_{{\mathcal{D}}^{\prime}}$, $p^{\prime}_{{\mathcal{D}}^{\prime}}$, and $p^{\prime\prime}_{{\mathcal{D}}^{\prime}}$ has, if it exists, in-degree at most one, and that there exists a directed path from each of $p_{{\mathcal{D}}^{\prime}}$, $p^{\prime}_{{\mathcal{D}}^{\prime}}$, and $p^{\prime\prime}_{{\mathcal{D}}^{\prime}}$ to a leaf in ${\mathcal{D}}^{\prime}$ that does not traverse a reticulation. Lastly, since ${\mathcal{D}}^{\prime}$ is tree-child, at most one of $u_{{\mathcal{D}}^{\prime}}$, $v_{{\mathcal{D}}^{\prime}}$, and $v^{\prime}_{{\mathcal{D}}^{\prime}}$ is a reticulation. Noting that $u_{{\mathcal{D}}^{\prime}}\neq\rho$, because $u\neq\rho$ by the definition of SNPR±, we next obtain a digraph ${\mathcal{D}}$ from ${\mathcal{D}}^{\prime}$ in one of the following five ways. The setup of ${\mathcal{D}}^{\prime}$ in each case is illustrated in Figure 5.
@@
[…]
Figure 5. The setup of ${\mathcal{D}}^{\prime}$ in all five cases (D1)–(D5) in the construction of ${\mathcal{D}}$ from ${\mathcal{D}}^{\prime}$ as detailed in the proof of Lemma 6.5. Dashed edges may or may not exist in ${\mathcal{D}}^{\prime}$.
@@
It now follows from (6.5.1) and (6.5.2) that ${\mathcal{D}}$ is an agreement tree-child digraph for ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$.
6.5.3.
@@
$|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}_{t}}|\leq|E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|+2.$
@@
[…]
Figure 6. An illustration of (R4) and (R6) as described in the proof of (6.5.3). Blue indicates edges and vertices of $R_{u}^{\prime}$ and $R_{u}$, and green indicates edges and vertices of $R_{c_{u}}^{\prime}$ and $R_{c_{u}}$.
@@
Refer to caption(R8)$v$$p_{u}$$u$$c_{u}$$v$$p_{u}$$c_{u}$${\mathcal{N}}_{t-1}$${\mathcal{N}}_{t}$$t_{1}$$t_{1}$$t$$s=s_{1}$$s=s_{1}$$t$$v$$p_{u}$$c_{u}$${\mathcal{N}}_{t}$$t_{1}$$t$$s=s_{1}$$s_{1}^{\prime}$$t_{1}^{\prime}$(R11)$v$$p_{u}$$u$$c_{u}$${\mathcal{N}}_{t-1}$$t_{1}$$t$$s=s_{1}$$s_{1}^{\prime}$$t_{1}^{\prime}$
Figure 7. An illustration of (R8) and (R11) as described in the proof of (6.5.3). Blue indicates edges and vertices of $R_{u}^{\prime}$ and $R_{u}$, and green indicates edges and vertices of $R_{c_{u}}^{\prime}$ and $R_{c_{u}}$.
@@
Now, let $C^{\prime}=E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}$, and let $C=E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}_{t}}$. To complete the proof of (6.5.3), we compare the number of edges in $C^{\prime}$ with the number of edges in $C$. First, observe that, if $(p_{u^{\prime}},c_{u^{\prime}})\in C^{\prime}$, then $(p_{u^{\prime}},u^{\prime})\in C$ and $(u^{\prime},c_{u^{\prime}})\notin C$. Furthermore, if $(p_{u^{\prime}},c_{u^{\prime}})\notin C^{\prime}$, then neither $(p_{u^{\prime}},u^{\prime})$ nor $(u^{\prime},c_{u^{\prime}})$ is in $C$. We next list the edges that are in $C^{\prime}$ but not in $C$ and vice versa for each of (R1)–(R11). While $C^{\prime}-C$ contains edges in ${\mathcal{N}}_{t-1}$ that are not edges in ${\mathcal{N}}_{t}$, the set $C-C^{\prime}$ contains edges in ${\mathcal{N}}_{t}$ that are not edges in ${\mathcal{N}}_{t-1}$. Thus, edges that are common to ${\mathcal{N}}_{t-1}$ and ${\mathcal{N}}_{t}$ and common to $C^{\prime}$ and $C$ are not considered in the following table. Moreover, regardless of which of (R1)–(R11) applies, we note that $C^{\prime}-C$ may or may not contain $(p_{u^{\prime}},c_{u^{\prime}})$ and $C-C^{\prime}$ may or may not contain $(p_{u^{\prime}},u^{\prime})$. However, $C^{\prime}-C$ contains $(p_{u^{\prime}},c_{u^{\prime}})$ if and only if $C-C^{\prime}$ contains $(p_{u^{\prime}},u^{\prime})$, and so we have also omitted in the table the possibility that $C^{\prime}-C$ may contain $(p_{u^{\prime}},c_{u^{\prime}})$ and the possibility that $C-C^{\prime}$ may contain $(p_{u^{\prime}},u^{\prime})$.
@@
Since $|C-C^{\prime}|\leq|C^{\prime}-C|+2$ in all cases, this completes the proof of (6.5.3). ∎
6.5.4.
@@
$|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}_{0}}|\leq|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+2.$
@@
Again, to ease reading, we view ${\mathcal{R}}^{\prime}_{0}$ simply as a collection of edges in ${\mathcal{N}}_{0}$ and describe the construction of ${\mathcal{R}}_{0}$ from ${\mathcal{R}}^{\prime}_{0}$ by edge deletions only. Let $P$ be the directed path in ${\mathcal{N}}_{0}$ that $e_{{\mathcal{D}}^{\prime}}$ corresponds to, and let $(s,t)$ be the first edge on $P$. Let $R_{s}^{\prime}$ be the element in ${\mathcal{R}}_{0}^{\prime}$ that contains $s$. We next construct an extension ${\mathcal{R}}_{0}$ of ${\mathcal{D}}$ in ${\mathcal{N}}_{0}$ by modifying $R_{s}^{\prime}$. This construction is similar to the constructions described in proof of (6.5.3), but much less involved. First, suppose that ${\mathcal{D}}$ has been obtained from ${\mathcal{D}}^{\prime}$ by applying (D1), (D2), or (D4). Then
@@
Now, let ${\mathcal{R}}_{0}=({\mathcal{R}}_{0}^{\prime}-\{R_{s}^{\prime}\})\cup\{R_{s}\}$. Since ${\mathcal{R}}_{0}^{\prime}$ is an extension of ${\mathcal{D}}^{\prime}$ in ${\mathcal{N}}_{0}$, a careful check shows that ${\mathcal{R}}_{0}$ is an extension of ${\mathcal{D}}$ in ${\mathcal{N}}_{0}$. Moreover, since each of (R1′′)–(R3′′) deletes at most two edges in obtaining $R_{s}$ from $R_{s}^{\prime}$, it follows that (6.5.4) holds. ∎
Finally, by combining Equation (1) with (6.5.1)and (6.5.4), we get
$\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$$\leq$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}_{0}}|+|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}_{t}}|)$
$\leq$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+2+|E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|+2)$
$\leq$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|)+\frac{1}{2}\cdot 4$
$\leq$ $w(\sigma_{1})+w(\sigma_{2})$
@@
Proof of Theorem 1.1. The theorem follows from Lemmas 6.4 and 6.5.∎
@@
The first equality follows from [6, Proposition 7.1] and the fact that each SNPR± contributes two to the weight of any SNPR sequence connecting ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$. Moreover, to establish the second equality, let $\sigma=({\mathcal{T}}={\mathcal{T}}_{0},{\mathcal{T}}_{1},{\mathcal{T}}_{2},\ldots,{\mathcal{T}}_{t}={\mathcal{T}}^{\prime})$ be an SNPR sequence connecting ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$. Then it follows from Lemma 6.4 and a careful inspection of the proof of Lemma 6.5 when applied to two phylogenetic $X$-trees that (D1) and, consequently, (R1)–(R4) and (R1) always apply. Hence, the last set of inequalities in the proof of Lemma 6.5 can be replaced with
$m_{\rm tc}({\mathcal{T}},{\mathcal{T}}^{\prime})$$\leq$ $|E_{{\mathcal{T}}_{0}}-E_{{\mathcal{R}}_{0}}|+|E_{{\mathcal{T}}_{t}}-E_{{\mathcal{R}}_{t}}|$
$\leq$ $|E_{{\mathcal{T}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+1+|E_{{\mathcal{T}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|+1$
$\leq$ $w(\sigma_{1})+w(\sigma_{2})$
@@
The next proposition shows that the bound given in Lemma 6.5 is essentially tight.
Proposition 6.7.

### v2

@@ 6 Bounding the tree-child SNPR distance
Hence, we may assume that $e\in{\mathcal{R}}^{\prime}_{t-1}$. Let $R^{\prime}_{u}$ be the element in ${\mathcal{R}}^{\prime}_{t-1}$ that contains $e$, let $R^{\prime}_{c_{u}}$ be the element in ${\mathcal{R}}^{\prime}_{t-1}$ that contains $c_{u}$, and let $R^{\prime}_{c_{u^{\prime}}}$ be the element in ${\mathcal{R}}^{\prime}_{t-1}$ that contains $c_{u^{\prime}}$. Recall that $R^{\prime}_{u}$, $R^{\prime}_{c_{u}}$, and $R^{\prime}_{c_{u^{\prime}}}$ are not necessarily pairwise distinct. If $(p_{u^{\prime}},c_{u^{\prime}})\in R^{\prime}_{c_{u^{\prime}}}$, then set $R_{c_{u^{\prime}}}$ to be the directed graph obtained from $R^{\prime}_{c_{u^{\prime}}}$ by subdividing $(p_{u^{\prime}},c_{u^{\prime}})$ with a new vertex $u^{\prime}$. Otherwise, if $(p_{u^{\prime}},c_{u^{\prime}})\notin R^{\prime}_{c_{u^{\prime}}}$, then set $R_{c_{u^{\prime}}}$ to be the directed graph obtained from $R^{\prime}_{c_{u^{\prime}}}$ by adding $(u^{\prime},c_{u^{\prime}})$. The construction is shown in Figure 10. Intuitively, $R_{c_{u^{\prime}}}$ is an extension of a component of a phylogenetic digraph in ${\mathcal{N}}_{t}$. Lastly, if $R_{u}^{\prime}=R^{\prime}_{c_{u^{\prime}}}$, then set $R^{\prime}_{u}=R_{c_{u^{\prime}}}$ and, if $R_{c_{u}}^{\prime}=R^{\prime}_{c_{u^{\prime}}}$, then set $R^{\prime}_{c_{u}}=R_{c_{u^{\prime}}}$ to account for the modification in obtaining $R_{c_{u^{\prime}}}$ from $R^{\prime}_{c_{u^{\prime}}}$.
@@
Figure 10: The two cases in the construction of $R_{c_{u^{\prime}}}$ from $R^{\prime}_{c_{u^{\prime}}}$. Blue indicates vertices and edges of $R_{c_{u^{\prime}}}$ and $R^{\prime}_{c_{u^{\prime}}}$.
@@
    If $W$ contains $u$ and $|W|\geqslant 2$, and $R^{\prime}_{u}\neq R^{\prime}_{c_{u}}$, then obtain $R_{u}$ from $R^{\prime}_{u}$ by deleting each vertex in $W$, and obtain $R_{c_{u}}$ from $R^{\prime}_{c_{u}}$ by adding $(p_{u},c_{u})$ and each edge of $R^{\prime}_{u}$ that joins two vertices in $W-\{u\}$.
@@
    If $W$ contains $u$ and $|W|\geqslant 2$, and $R^{\prime}_{u}=R^{\prime}_{c_{u}}$, then obtain $R_{u}$ from $R^{\prime}_{u}$ by deleting $u$ and adding $(p_{u},c_{u})$.
As an aside, recall that we are dealing with extensions (and not with the more restricted root extensions). Indeed, if $c_{u}$ is a reticulation in ${\mathcal{N}}_{t-1}$ and a vertex with in-degree one and out-degree one in $R_{c_{u}}^{\prime}$, then $R_{c_{u}}$ is an extension and not a root extension. Now, regardless of which of (R1)–(R4) applies, let
${\mathcal{R}}^{\prime}_{t}=({\mathcal{R}}^{\prime}_{t-1}-\{R^{\prime}_{u},R^{\prime}_{c_{u}},R^{\prime}_{c_{u^{\prime}}}\})\cup\{R_{u},R_{c_{u}},R_{c_{u^{\prime}}}\}.$
It is easily checked that ${\mathcal{R}}^{\prime}_{t}$ is an extension of $\mathcal{D}$$^{\prime}$ in ${\mathcal{N}}_{t}$ with
@@
$\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ $\leqslant$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}^{\prime}_{t}}|)$
@@
We complete the proof of the lemma by assuming that $e\in{\mathcal{M}}^{\prime}_{t-1}$. Let $e_{{\mathcal{D}}^{\prime}}=(u_{{\mathcal{D}}^{\prime}},v_{{\mathcal{D}}^{\prime}})$ be the unique edge in $\mathcal{D}$$^{\prime}$ that $e$ corresponds to. If $u_{{\mathcal{D}}^{\prime}}$ has in-degree two in $\mathcal{D}$$^{\prime}$, let $p_{{\mathcal{D}}^{\prime}}$ and $p^{\prime}_{{\mathcal{D}}^{\prime}}$ be the two parents of $u_{{\mathcal{D}}^{\prime}}$. Furthermore, if $u_{{\mathcal{D}}^{\prime}}$ has out-degree two, let $v^{\prime}_{{\mathcal{D}}^{\prime}}$ be the child of $u_{{\mathcal{D}}^{\prime}}$ that is not $v_{{\mathcal{D}}^{\prime}}$ and, if $v^{\prime}_{{\mathcal{D}}^{\prime}}$ is a reticulation, let $p^{\prime\prime}_{{\mathcal{D}}^{\prime}}$ be the parent of $v^{\prime}_{{\mathcal{D}}^{\prime}}$ that is not $u_{{\mathcal{D}}^{\prime}}$. Since $\mathcal{D}$$^{\prime}$ is tree-child, observe that each of $p_{{\mathcal{D}}^{\prime}}$, $p^{\prime}_{{\mathcal{D}}^{\prime}}$, and $p^{\prime\prime}_{{\mathcal{D}}^{\prime}}$ has, if it exists, in-degree at most one, and that there exists a directed path from each of $p_{{\mathcal{D}}^{\prime}}$, $p^{\prime}_{{\mathcal{D}}^{\prime}}$, and $p^{\prime\prime}_{{\mathcal{D}}^{\prime}}$ to a leaf in $\mathcal{D}$$^{\prime}$ that does not traverse a reticulation. Lastly, since $\mathcal{D}$$^{\prime}$ is tree-child, at most one of $u_{{\mathcal{D}}^{\prime}}$, $v_{{\mathcal{D}}^{\prime}}$, and $v^{\prime}_{{\mathcal{D}}^{\prime}}$ is a reticulation. Noting that $u_{{\mathcal{D}}^{\prime}}\neq\rho$, because $u\neq\rho$ by the definition of SNPR±, we next obtain a digraph $\mathcal{D}$} from $\mathcal{D}$$^{\prime}$ in one of the following five ways, which are illustrated in Figure 11.
[…]
Figure 11: The setup of $\mathcal{D}$$^{\prime}$ and $\mathcal{D}$} in all five cases (D1)–(D5) in the construction of $\mathcal{D}$} from $\mathcal{D}$$^{\prime}$ as detailed in the proof of Lemma 14. Dashed edges may or may not exist in $\mathcal{D}$$^{\prime}$ and $\mathcal{D}$}. Vertex labels $y_{1}$, $y_{2}$, and $y_{3}$ are only used to clarify the figure.
@@
It now follows from (14.1) and (14.2) that $\mathcal{D}$} is an agreement tree-child digraph for ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$.
14.3.
@@
$|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}_{t}}|\leqslant|E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|+2.$
@@
[…]
Figure 12: An illustration of (R4) and (R6) as described in the proof of (14.3). Blue indicates edges and vertices of $R_{u}^{\prime}$ and $R_{u}$, and green indicates edges and vertices of $R_{c_{u}}^{\prime}$ and $R_{c_{u}}$.
@@
[…]
Figure 13: An illustration of (R8) and (R11) as described in the proof of (14.3). Blue indicates edges and vertices of $R_{u}^{\prime}$ and $R_{u}$, and green indicates edges and vertices of $R_{c_{u}}^{\prime}$ and $R_{c_{u}}$.
@@
Now, let $C^{\prime}=E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}$, and let $C=E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}_{t}}$. To complete the proof of (14.3), we compare the number of edges in $C^{\prime}$ with the number of edges in $C$. First, observe that, if $(p_{u^{\prime}},c_{u^{\prime}})\in C^{\prime}$, then $(p_{u^{\prime}},u^{\prime})\in C$ and $(u^{\prime},c_{u^{\prime}})\notin C$. Furthermore, if $(p_{u^{\prime}},c_{u^{\prime}})\notin C^{\prime}$, then neither $(p_{u^{\prime}},u^{\prime})$ nor $(u^{\prime},c_{u^{\prime}})$ is in $C$. We next list the edges that are in $C^{\prime}$ but not in $C$ and vice versa for each of (R1)–(R11). While $C^{\prime}-C$ contains edges in ${\mathcal{N}}_{t-1}$ that are not edges in ${\mathcal{N}}_{t}$, the set $C-C^{\prime}$ contains edges in ${\mathcal{N}}_{t}$ that are not edges in ${\mathcal{N}}_{t-1}$. Thus, edges that are common to ${\mathcal{N}}_{t-1}$ and ${\mathcal{N}}_{t}$ and common to $C^{\prime}$ and $C$ are not considered in the following table. Moreover, regardless of which of (R1)–(R11) applies, we note that $C^{\prime}-C$ may or may not contain $(p_{u^{\prime}},c_{u^{\prime}})$ and $C-C^{\prime}$ may or may not contain $(p_{u^{\prime}},u^{\prime})$. However, $C^{\prime}-C$ contains $(p_{u^{\prime}},c_{u^{\prime}})$ if and only if $C-C^{\prime}$ contains $(p_{u^{\prime}},u^{\prime})$, and so we have also omitted in the table the possibility that $C^{\prime}-C$ may contain $(p_{u^{\prime}},c_{u^{\prime}})$ and the possibility that $C-C^{\prime}$ may contain $(p_{u^{\prime}},u^{\prime})$.
@@
Since $|C-C^{\prime}|\leqslant|C^{\prime}-C|+2$ in all cases, this completes the proof of (14.3). ∎
14.4.
@@
$|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}_{0}}|\leqslant|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+2.$
@@
Again, to ease reading, we view ${\mathcal{R}}^{\prime}_{0}$ simply as a collection of edges in ${\mathcal{N}}_{0}$ and describe the construction of ${\mathcal{R}}_{0}$ from ${\mathcal{R}}^{\prime}_{0}$ by edge deletions only. Let $P$ be the directed path in ${\mathcal{N}}_{0}$ that $e_{{\mathcal{D}}^{\prime}}$ corresponds to, and let $(s,t)$ be the first edge on $P$. Let $R_{s}^{\prime}$ be the element in ${\mathcal{R}}_{0}^{\prime}$ that contains $s$. We next construct an extension ${\mathcal{R}}_{0}$ of $\mathcal{D}$} in ${\mathcal{N}}_{0}$ by modifying $R_{s}^{\prime}$. This construction is similar to the constructions described in proof of (14.3), but much less involved. First, suppose that $\mathcal{D}$} has been obtained from $\mathcal{D}$$^{\prime}$ by applying (D1), (D2), or (D4). Then
@@
Now, let ${\mathcal{R}}_{0}=({\mathcal{R}}_{0}^{\prime}-\{R_{s}^{\prime}\})\cup\{R_{s}\}$. Since ${\mathcal{R}}_{0}^{\prime}$ is an extension of $\mathcal{D}$$^{\prime}$ in ${\mathcal{N}}_{0}$, a careful check shows that ${\mathcal{R}}_{0}$ is an extension of $\mathcal{D}$} in ${\mathcal{N}}_{0}$. Moreover, since each of (R1′′)–(R3′′) deletes at most two edges in obtaining $R_{s}$ from $R_{s}^{\prime}$, it follows that (14.4) holds. ∎
Finally, by combining Equation (1) with (14.1)and (14.4), we get
$\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ $\leqslant$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}_{0}}|+|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}_{t}}|)$
$\leqslant$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+2+|E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|+2)$
$\leqslant$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|)+\frac{1}{2}\cdot 4$
$\leqslant$ $w(\sigma_{1})+w(\sigma_{2})$
@@
Proof of Theorem 1. The theorem follows from Lemmas 13 and 14.∎
@@
The first equality follows from Bordewich et al. [7, Proposition 7.1] and the fact that each SNPR± contributes two to the weight of any SNPR sequence connecting $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$. Moreover, to establish the second equality, let
$\sigma=({\mathcal{T}}={\mathcal{T}}_{0},{\mathcal{T}}_{1},{\mathcal{T}}_{2},\ldots,{\mathcal{T}}_{t}={\mathcal{T}}^{\prime})$
be an SNPR sequence connecting $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$. Then it follows from Lemma 13 and a careful inspection of the proof of Lemma 14 when applied to two phylogenetic $X$-trees that (D1) and, consequently, (R1)–(R4) and (R1) always apply. Hence, the last set of inequalities in the proof of Lemma 14 can be replaced with
$m_{\rm tc}({\mathcal{T}},{\mathcal{T}}^{\prime})$ $\leqslant$ $|E_{{\mathcal{T}}_{0}}-E_{{\mathcal{R}}_{0}}|+|E_{{\mathcal{T}}_{t}}-E_{{\mathcal{R}}_{t}}|$
$\leqslant$ $|E_{{\mathcal{T}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+1+|E_{{\mathcal{T}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|+1$
$\leqslant$ $w(\sigma_{1})+w(\sigma_{2})$
@@
The next proposition shows that the bound given in Lemma 14 is essentially tight.
[…]
Figure 14: Two phylogenetic networks $\mathcal{N}$$_{\ell}$ and $\mathcal{N}$$_{\ell}^{\prime}$ on $n=3\ell$ leaves with $\ell>1$, an agreement tree-child digraph $\mathcal{D}$$_{\ell}$ for $\mathcal{N}$$_{\ell}$ and $\mathcal{N}$$_{\ell}^{\prime}$, an extension of $\mathcal{D}$$_{\ell}$ in $\mathcal{N}$$_{\ell}$ and an extension of $\mathcal{D}$$_{\ell}$ in $\mathcal{N}$$_{\ell}^{\prime}$ indicated in red. This example shows that the bound given in Lemma 14 is essentially tight. For details, see the proof of Proposition 16.
Proposition 16.